Given the following information relating to the yields to maturity on several one- year, zero-coupon bonds:
Bond Yield (%)
US Treasury 3.0
AAA Corporate 3.3
A Corporate 3.9
BB Corporate 4.8
a. Find the price of a one-year, zero-coupon corporate bond with a AAA rating.
b. Find the credit spread on the AAA-rated corporate bonds.
c. Find the credit spread on the A-rated corporate bonds.
d. Find the credit spread on the BB-rated corporate bonds.
e. In what way does the credit spread change with the bond rating?
1
Expert's answer
2019-12-03T10:11:50-0500

a. The price of a one-year, zero-coupon corporate bond with a AAA rating is: P = FV/(1 + 0.033), where FV is face value.

b. The credit spread on the AAA-rated corporate bonds is: 3.3 - 3.0 = 0.3%.

c. The credit spread on the A-rated corporate bonds is: 3.9 - 3.0 = 0.9%.

d. The credit spread on the BB-rated corporate bonds is: 4.8 - 3.0 = 1.8%.

e. The credit spread increases with the decrease in bond rating.